# RetinexDualV2: Physically-Grounded Dual Retinex for Generalized UHD Image Restoration

### 30 Mar 2026 v1

@@ Abstract
We propose RetinexDualV2, a unified, physically grounded dual-branch framework for diverse Ultra-High-Definition (UHD) image restoration. Unlike generic models, our method employs a Task-Specific Physical Grounding Module (TS-PGM) to extract degradation-aware priors (e.g., rain masks and dark channels). These explicitly guide a Retinex decomposition network via a novel Physical-conditioned Multi-head Self attention (PC-MSA) mechanism, enabling robust reflection and illumination correction. This physical conditioning allows a single architecture to handle various complex degradations seamlessly, without task-specific structural modifications. RetinexDualV2 demonstrates exceptional generalizability, securing 4th place in the NTIRE 2026 Day and Night Raindrop Removal Challenge and 5th place in the Joint Noise Low-light Enhancement (JNLLIE) Challenge. Extensive experiments confirm the state-of-the-art performance and efficiency of our physically motivated approach.
@@ 1 Introduction
Although imaging sensors and displays have advanced significantly to deliver highly detailed and sophisticated Ultra-High-Definition (UHD) images, restoring captured information distorted or degraded by light exposure, haze, shadow, or weather conditions has grown increasingly difficult. This causes a bottleneck for more advanced tasks, such as object detection, semantic segmentation, and visual reasoning. Accordingly, images restoration (IR) became more crucial not only for the plausibility of the image but also for the improvement of computer vision in general.
Numerous learning-based methods have been surfaced to tackle image restoration challenges [17, 28, 2], substantially improving the removal of various degradation artifacts. Nevertheless, these techniques struggle to process UHD images effectively, as the enormous data volume and resolution demand excessive computational resources. This constraint has become a major obstacle in achieving high-quality UHD image restoration.
@@
Alternatively, Retinexformer [2] introduced a Retinex-based model for low-light enhancement, inspiring several follow-up works [1, 15]. Although these methods achieved top performance in low-light restoration, they generally treat reflectance and illumination as homogeneous components and do not extend Retinex theory into a more general framework for broader image restoration tasks beyond illumination/color correction. RetinexDual [11] addressed this problem by proposing an approach that recongizes the duel nature of retinex decomposition. However, it faces challenges in certain complex scenarios as there is no physical-grounding for the learning process.
@@
Based on this, we propose RetinexDualV2, a physically grounded evolution of RetinexDual that preserves its effective dual-branch handling of reflectance and illumination while infusing task-specific physical priors to guide restoration. Specifically, the degraded input is fed into two parallel extractors: a Retinex decomposer and a task-specific physical grounding module. The resulting reflectance and illumination components are then processed through Physical-Grounded Scale Attentive Mamba (PG-SAMBA) for spatially adaptive, detail-preserving reflectance correction and Physical-Grounded Frequency Illumination Adaptor (PG-FIA) for physically consistent illumination adjustment. Within each layer, the physical features are incorporated using Physical-conditioned Multi-head Self attention (PC-MSA). The main contributions are summerized as follows:
@@
  • In PC-MSA, we inject the physical map into the restoration process using modulated self attention mechanism to regularize outputs and ensure physical plausibility.
@@ 2 Related Work
Ultra-High-Definition Image Restoration. With the growing demand for high-fidelity visual content, extensive research has been dedicated to Ultra-High-Definition (UHD) image restoration [18, 33, 34, 20]. For instance, LLFormer [20] relies on fusion cross-attention to merge features across different stages, alongside axis-based multi-head self-attention to manage multi-scale representations. While highly effective at retaining UHD details, this architecture demands massive computational resources, making it impractical for standard hardware. Alternatively, methods like Wave-Mamba [34] tackle downsampling-related information loss by processing distinct frequency bands in parallel through specialized Mamba blocks. Similarly, ERR [33] introduces a three-stage pipeline specifically designed to handle zero-, low-, and high-frequency components separately. Even though these frequency-domain techniques successfully mitigate computational bottlenecks and maintain fine details, they are not a universal solution. Specifically, localized spatial degradations are typically more effectively addressed directly within the spatial domain, as this approach provides a more precise representation of the mapping between corrupted and pristine image regions [11].
Retinex Theory in Image Restoration. Low-Light Image Enhancement (LLIE) has recently shifted focus toward Retinex theory, leveraging the idea that image noise is primarily concentrated in the illumination layer [2, 1, 15]. Retinexformer [2] successfully applied this concept via an illumination-guided transformer, inspiring a wave of state-of-the-art extensions including RetinexMamba [1], Retinexformer+ [15], ERetinex [9], and MambaLLIE [24]. Despite their strong performance, these models exhibit structural shortcomings: they either modify only one Retinex component or process both using the exact same architectural design, and they largely ignore Retinex principles for broader Image Restoration (IR) applications. RetinexDual [11] tackled this issue directly by formulating an architecture tailored to the dual nature of Retinex components. However, this method encounters difficulties in specialized cases, demonstrating that relying solely on Retinex theory is inadequate for universal restoration. To overcome these limitations and effectively process highly complex scenes, incorporating explicit physical grounding is essential.
Physical Grounding in Image Restoration. Physical priors derived from the optical properties of image formation serve as foundational regularizers for ill-posed restoration tasks. By explicitly modeling the underlying degradation mechanisms ,such as leveraging the dark channel prior for atmospheric scattering [6] or utilizing log-chromaticity and shadow-resistant representations for illumination invariance [21]. These approaches effectively isolate the clean signal from complex perturbations. Furthermore, incorporating explicit structural priors, such as structure-guided designs in scale-aware low-light enhancement [7], explicitly utilizes physical geometry to constrain the restoration space. Integrating these explicit physical and structural constraints into modern frameworks enables models to maintain strict radiometric consistency and robust edge preservation across a broad spectrum of adverse conditions, from extreme low-light environments to weather-induced artifacts. Consequently, this guides the learning process toward physically plausible image recovery, proving highly resilient where purely data-driven statistical mappings often fail.
@@ 3 Methodology
Our key contribution in this paper is the design of a task-specific physical grounding module (TS-PGM) 3.1 that can be adjusted to the nature of degradation while integrating this physical prior with the dual branch of retinex decomposition using a Physical-conditioned Multi-head Self attention (PC-MSA) in the RetinxDualV2 model 3.2. The proposed RetinexDualV2 is presented in Fig. 3, which contains two sub-networks, PG-SAMBA and PG-FIA, retinex decomposer, and a task-based physical grounding module. In each of the branches, Physical-Grounded Scale Adaptive Mamba Block (PG-SAMB) and Physical-Grounded Fourier Correction Block (PG-FCB) are responsible for restoring the reflectance and illumination components, respectively, while relying on the physical conditioning provided by (TS-PGM).
@@
Figure 3: RetinexDualV2 Overview. Based on retinex theory, it decomposes the UHD image into $R_{eff}$ and $L_{eff}$ and operates on them using 2 sub-networks: Physical-Grounded Scale Attentive MaMBA (PG-SAMBA), and Physical-Grounded Frequency Illumination Adaptor (PG-FIA), respectively. Both branches recieves guidance $p_{m}$ from the Task-Specific Physical Grounding Module (TS-PGM). Noting that each output from a different level has a convolution layer before it for processing, which wasn’t illustrated for simplicity.
@@ 3.1 Task-Specific Physical Grounding Module (TS-PGM)
The TS-PGM module extracts task-dependent physical priors as shown in Fig. 2, guiding the restoration process by modulating feature interactions through PC-MSA. Unlike general-purpose feature extraction, TS-PGM explicitly encodes degradation-specific physics to condition both reflection and illumination branches. Across all tasks, the physical prior is computed by fusing gradient and intensity-based representations:
@@
$=\mathbf{G}_{\text{cm}}\cdot I,$ (4)
@@
3.2 RetinexDualV2 Framework:
Inspired from RetinexDual [11], RetinexDualV2 follows Retinex theory, which decomposes a distorted image $I\in\mathbb{R}^{H\times W\times 3}$ as:
@@ 3.2 RetinexDualV2 Framework:
Physical-conditioned Multi-head Self attention (PC-MSA) The core component enabling physical grounding across both branches is the Physical-conditioned Multi-head Self attention (PC-MSA). PC-MSA extends standard multi-head self-attention by modulating feature interactions with task-specific physical guidance extracted from TS-PGM.
@@ 4 Experiments
Figure 4: Visual comparison between different architectures on UHD-$L$ [14].
@@
Figure 5: Visual comparison between different architectures on 4K-Rain13K dataset [5].
Table 3: Final ranking of NTIRE 2026 Day and Night Raindrop Removal for Dual-Focused Images Challenge.
@@
Table 4: Final ranking of NTIRE 2026 Joint Noise Low-light enhancement Challenge.
@@ 4.1 Experiment Settings
Datasets. Our model was evaluated on 4K-Rain13K dataset [5] for deraining, and UHD-LL dataset [14] for low-light enhancement. Additionally, we participated in the NTIRE 2026 Challenges on Day and Night Raindrop Removal for Dual-Focused Images, Joint Noise Low-light enhancement, Nighttime Dehazing, and Shadow Removal, where our method was evaluated on the respective challenge datasets.
Implementation details. RetinexDuelV2 model was trained on using Nvidea H100 GPU with the following settings. Using AdamW optimizer [16] with an initial learning rate of 2e-5, which is gradually reduced to 1e-7 through cosine annealing learning rate scheduler. The patch size used is 768×768 with a batch size of 5.
@@ 4.2 Quantitative and Qualitative Results
Dehazing
@@
Performance on Competitions. Our proposed RetinexDualV2 demonstrates strong generalization capabilities across multiple image restoration tasks in the NTIRE 2026 Challenges. Specifically, our model secured the 4th place in the Day and Night Raindrop Removal for Dual-Focused Images Challenge (Table 3), achieving a PSNR of 27.24 dB, SSIM of 0.806, and LPIPS of 0.289. In the Joint Noise Low-light Enhancement Challenge (Table 4), we achieved the 5th place with a PSNR of 18.69 dB, SSIM of 0.65, and LPIPS of 0.50. Furthermore, the model achieved competitive performance in the Nighttime Dehazing (PSNR 24.87 dB, SSIM 0.868, LPIPS 0.134) and Shadow Removal (PSNR 23.09 dB, SSIM 0.823, LPIPS 0.107) competitions. This consistent performance across diverse and challenging datasets highlights the robustness and adaptability of the Retinex-based dual architecture. It proves that our unified framework can be effectively generalized to various complex, real-world image restoration tasks without requiring task-specific architectural modifications, all while maintaining an efficient model complexity. Also, the visual results in Figure 6 demonstrate the qualitative improvements achieved by our method across different restoration tasks, showcasing clearer details and reduced artifacts in the enhanced images.
@@
Ablation Study. To dissect the effectiveness of our physical-guided structural designs and the efficacy of the illumination-reflection decomposition, we conducted an ablation study on the UHD-LL dataset [14]. Specifically, we focus on three core aspects: the dual-branch Retinex decomposition architecture, the extraction of task-specific physical priors via the Task-Specific Physical Grounding Module (TS-PGM), and their targeted injection through the Physical-conditioned Multi-head Self attention (PC-MSA). As reported in Table 5, utilizing a single-branch architecture (PG-SAMBA) without decoupling illumination and reflection (“w/o Dual-Branch”) yields the lowest performance at 26.67 dB PSNR and 0.919 SSIM. Introducing the dual-branch design but ignoring the synthesized physical prior entirely (“w/o Physical Prior”) provides a significant improvement up to 28.79 dB PSNR and 0.920 SSIM. If the network extracts the physical prior but relies on standard feature interactions by concatnating the prior instead of selectively conditioning via PC-MSA (“w/o PC-MSA”), it achieves 28.43 dB PSNR and 0.927 SSIM. Finally, our full RetinexDualV2 model achieves the optimal performance of 28.91 dB PSNR and 0.935 SSIM, demonstrating that dynamically synthesizing task-specific physical priors and explicitly steering the dual-branch feature representation via PC-MSA are strictly necessary for maximizing restoration capabilities.

### 09 Apr 2026 v2

@@ Abstract
We propose RetinexDualV2, a unified, physically grounded dual-branch framework for diverse Ultra-High-Definition (UHD) image restoration. Unlike generic models, our method employs a Task-Specific Physical Grounding Module (TS-PGM) to extract degradation-aware priors (e.g., rain masks and dark channels). These explicitly guide a Retinex decomposition network via a novel Physical-Conditioned Multi-head Self-Attention (PC-MSA) mechanism, enabling robust reflection and illumination correction. This physical conditioning allows a single architecture to handle various complex degradations seamlessly, without task-specific structural modifications. RetinexDualV2 demonstrates exceptional generalizability, securing 4th place in the NTIRE 2026 Day and Night Raindrop Removal Challenge and 5th place in the Joint Noise Low-light Enhancement (JNLLIE) Challenge. Extensive experiments confirm the state-of-the-art performance and efficiency of our physically motivated approach. Code is available at https://github.com/ErrorLogic1211/RetinexDual/tree/master/RetinexDualV2
@@ 1 Introduction
Although imaging sensors and displays have advanced significantly to deliver highly detailed and sophisticated Ultra-High-Definition (UHD) images, restoring captured information distorted or degraded by light exposure, haze, shadow, or weather conditions has grown increasingly difficult. This causes a bottleneck for more advanced tasks, such as object detection, semantic segmentation, and visual reasoning. Accordingly, image restoration (IR) became more crucial not only for the plausibility of the image but also for the improvement of computer vision in general.
Numerous learning-based methods have emerged to tackle image restoration challenges [20, 32, 3], substantially improving the removal of various degradation artifacts. Nevertheless, these techniques struggle to process UHD images effectively, as the enormous data volume and resolution demand excessive computational resources. This constraint has become a major obstacle in achieving high-quality UHD image restoration.
@@
Alternatively, Retinexformer [3] introduced a Retinex-based model for low-light enhancement, inspiring several follow-up works [2, 18]. Although these methods achieved top performance in low-light restoration, they generally treat reflectance and illumination as homogeneous components and do not extend Retinex theory into a more general framework for broader image restoration tasks beyond illumination/color correction. RetinexDual [13] addressed this problem by proposing an approach that recognizes the dual nature of retinex decomposition. However, it faces challenges in certain complex scenarios as there is no physical-grounding for the learning process.
@@
Based on this, we propose RetinexDualV2, a physically grounded evolution of RetinexDual that preserves its effective dual-branch handling of reflectance and illumination while infusing task-specific physical priors to guide restoration. Specifically, the degraded input is fed into two parallel extractors: a Retinex decomposer and a task-specific physical grounding module. The resulting reflectance and illumination components are then processed through Physical-Grounded Scale Attentive Mamba (PG-SAMBA) for spatially adaptive, detail-preserving reflectance correction and Physical-Grounded Frequency Illumination Adaptor (PG-FIA) for physically consistent illumination adjustment. Within each layer, the physical features are incorporated using Physical-Conditioned Multi-head Self-Attention (PC-MSA). The main contributions are summarized as follows:
@@
  • In PC-MSA, we inject the physical map into the restoration process using a modulated self-attention mechanism to regularize outputs and ensure physical plausibility.
@@ 2 Related Work
Ultra-High-Definition Image Restoration. Extensive research addresses UHD restoration [22, 37, 38, 24]. Approaches like LLFormer [24] retain UHD details via multi-scale attention but demand massive computational resources. Alternatively, frequency-domain methods like Wave-Mamba [38] and ERR [37] mitigate computational bottlenecks by processing distinct spectral bands. However, they lack universal applicability; localized spatial degradations often require direct spatial-domain processing to accurately map corrupted regions to pristine counterparts [13].
Retinex Theory in Image Restoration. Recent Low-Light Image Enhancement (LLIE) increasingly adopts Retinex theory, treating illumination as the primary noise carrier [3, 2, 18]. While Retinexformer [3] and its extensions [2, 18, 11, 28] exhibit strong performance, they typically process illumination and reflectance homogeneously and neglect broader image restoration applications. RetinexDual [13] introduced a dual-branch architecture tailored to each component, yet it struggles in specialized scenarios, indicating that pure Retinex theory requires explicit physical grounding for reliable universal restoration.
Physical Grounding in Image Restoration. Physical priors provide essential regularization for ill-posed restoration tasks. Methods explicitly modeling degradation mechanism, such as the dark channel prior for atmospheric scattering [7] or log-chromaticity for illumination invariance [25], effectively isolate clean signals. Furthermore, structural priors [9] geometrically constrain the restoration space. Integrating these explicit priors ensures strict radiometric consistency and robust edge preservation, guiding models toward physically plausible, resilient recovery where purely data-driven statistical mappings often fail.
@@ 3 Methodology
Our key contribution in this paper is the design of a task-specific physical grounding module (TS-PGM) 3.1 that can be adjusted to the nature of degradation while integrating this physical prior with the dual branch of retinex decomposition using a Physical-conditioned Multi-head Self-Attention (PC-MSA) in the RetinexDualV2 model 3.2. The proposed RetinexDualV2 is presented in Fig. 3, which contains two sub-networks, PG-SAMBA and PG-FIA, retinex decomposer, and a task-based physical grounding module. In each of the branches, Physical-Grounded Scale Adaptive Mamba Block (PG-SAMB) and Physical-Grounded Fourier Correction Block (PG-FCB) are responsible for restoring the reflectance and illumination components, respectively, while relying on the physical conditioning provided by (TS-PGM).
@@
Figure 3: RetinexDualV2 Overview. Based on retinex theory, it decomposes the UHD image into $R_{eff}$ and $L_{eff}$ and operates on them using 2 sub-networks: Physical-Grounded Scale Attentive MaMBA (PG-SAMBA), and Physical-Grounded Frequency Illumination Adaptor (PG-FIA), respectively. Both branches receive guidance $p_{m}$ from the Task-Specific Physical Grounding Module (TS-PGM). Noting that each output from a different level has a convolution layer before it for processing, which wasn’t illustrated for simplicity.
@@ 3.1 Task-Specific Physical Grounding Module (TS-PGM)
The TS-PGM module extracts task-dependent physical priors as shown in Fig. 2, guiding the restoration process using either degradation resistant components (as in shadow prior) or degradation sensitive components (as in rain prior) by modulating feature interactions through PC-MSA. Unlike general-purpose feature extraction, TS-PGM explicitly encodes degradation-specific physics to condition both reflection and illumination branches. Across all tasks, the physical prior is computed by fusing gradient and intensity-based representations:
@@
$[E,E_{l},E_{ll}]$ $=\mathbf{G}_{\text{cm}}\cdot I,$ (4)
@@
Motivation for Degradation-Specific Priors. By extracting task-dependent structural representations rather than generic convolutional features, TS-PGM serves as a lightweight task-selector. The overarching motivation for utilizing distinct priors (rain mask, structure-aware illumination, dark channel, and log-chromaticity) stems from the observation that different degradations violate Retinex assumptions in fundamentally different ways. For instance, rain constitutes high-frequency localized occlusions violating reflectance continuity, whereas haze globally distorts illumination transmission. By injecting these carefully chosen physical mappings dynamically into PC-MSA, our model gracefully generalizes its single dual-branch backbone across drastically different conditions without requiring any task-specific rewriting of the core components.
3.2 RetinexDualV2 Framework
Inspired by RetinexDual [13], RetinexDualV2 follows Retinex theory, which decomposes a distorted image $I\in\mathbb{R}^{H\times W\times 3}$ as:
@@ 3.2 RetinexDualV2 Framework
Physical-Conditioned Multi-head Self-Attention (PC-MSA) The core component enabling physical grounding across both branches is the Physical-conditioned Multi-head Self-Attention (PC-MSA). PC-MSA extends standard multi-head self-attention by modulating feature interactions with task-specific physical guidance extracted from TS-PGM.
@@ 4 Experiments
Figure 4: Visual comparison between different architectures on UHD-$L$ [16]. Zoomed-in patches highlight our method’s superiority in noise suppression and reliable structure preservation in comparison with literature methods.
@@
Figure 5: Visual comparison between different architectures on 4K-Rain13K dataset [6]. Zoomed-in regions display that RetinexDualV2 removes localized rain artifacts optimally compared against recent literature.
Table 3: Final ranking of NTIRE 2026 The Second Challenge on Day and Night Raindrop Removal for Dual-Focused Images [17].
@@
Table 4: Final ranking of NTIRE 2026 Joint Noise Low-light enhancement Challenge [8].
@@ 4.1 Experiment Settings
Datasets. Our model was evaluated on 4K-Rain13K dataset [6] for deraining, and UHD-LL dataset [16] for low-light enhancement. Additionally, we participated in NTIRE 2026 The Second Challenge on Day and Night Raindrop Removal for Dual-Focused Images, Joint Noise Low-light enhancement, Nighttime Dehazing, and Shadow Removal, where our method was evaluated on the respective challenge datasets.
Implementation details. RetinexDualV2 model was trained using an NVIDIA H100 GPU with the following settings. Using AdamW optimizer [19] with an initial learning rate of 2e-5, which is gradually reduced to 1e-7 through a cosine annealing learning rate scheduler. The patch size used is 768×768 with a batch size of 5.
@@ 4.2 Quantitative and Qualitative Results
Dehazing Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Performance on Competitions. Our unified RetinexDualV2 model demonstrates robust generalization across multiple NTIRE 2026 Challenge restoration tasks without task-specific modifications or compromising efficiency. It secured 4th place in NTIRE 2026 The Second Challenge on Day and Night Raindrop Removal for Dual-Focused Images [17] (Table 3) with PSNR 27.24 dB, SSIM 0.806, and LPIPS 0.289, and 5th place in Joint Noise Low-light Enhancement [8] (Table 4) achieving PSNR 18.69 dB, SSIM 0.65, and LPIPS 0.50. It also performed competitively in Nighttime Dehazing [1] (PSNR 24.87 dB, SSIM 0.868, LPIPS 0.134) and Shadow Removal [21] (PSNR 23.09 dB, SSIM 0.823, LPIPS 0.107). Qualitatively, Figure 6 shows our method yields clearer details and reduced artifacts across these diverse datasets.
Complexity vs. Performance Trade-off. While RetinexDualV2 introduces a marginal parameter overhead ($\sim$0.12M, roughly a 2.5% increase) compared to its predecessor RetinexDual, this addition encapsulates the dedicated TS-PGM and PC-MSA modules. This trade-off is intentionally designed: rather than chasing unbounded parameter scaling for pure metric inflation, this slight capacity increase effectively grounds the architecture in solid physical principles. It yields subtle but critical quantitative improvements (e.g., +0.06 dB on 4K-Rain13K, +0.05 dB on UHD-LL) while substantially improving the model’s interpretability and stability across diverse, complex, and localized degradation scenarios, benefits that blindly scaling a data-driven model cannot reliably guarantee.
@@
Table 6: Ablation study comparing different fusion strategies on the validation set of NTIRE 2026 The Second Challenge on Day and Night Raindrop Removal for Dual-Focused Images.
[…]
W/o Fusion 27.11 0.818 0.256
MinMask 26.90 0.809 0.230
Median 27.57 0.826 0.251
Softblend (Ours) 27.70 0.829 0.244
Ablation Study. We conduct ablation studies to validate the physical grounding components and the rain removal fusion strategy in our pipeline.
1) Physical Grounding Components. We first illustrate the variations of our architectural components evaluated on the UHD-LL dataset [16] (Table 5): (i) w/o Dual-Branch: a single-branch configuration (PG-SAMBA) without illumination-reflection decoupling. (ii) w/o Physical Prior: utilizes the dual-branch architecture but discards structural priors entirely. (iii) w/o PC-MSA: extracts priors via TS-PGM but relies on simple concatenation rather than selective conditioning. (iv) Full Model: explicitly steers representations via PC-MSA. In terms of performance, the w/o Dual-Branch configuration yields only 26.67 dB PSNR and 0.919 SSIM. Incorporating the dual-branch structure (w/o Physical Prior) increases performance to 28.79 dB PSNR and 0.920 SSIM. Relying on simple concatenation (w/o PC-MSA) achieves 28.43 dB PSNR and 0.927 SSIM. Finally, our full RetinexDualV2 model reaches an optimal 28.91 dB PSNR and 0.935 SSIM, confirming the necessity of these targeted physical designs.
2) Fusion Strategy for Rain Removal. We evaluate four fusion strategies on the validation set of NTIRE 2026 The Second Challenge on Day and Night Raindrop Removal for Dual-Focused Images (Table 6): (i) W/o Fusion: a baseline without any fusion. (ii) MinMask: per-pixel Minimum Mask fusion. (iii) Median: takes the median fusion across scene predictions. (iv) Softblend (Ours): weights predictions inversely proportional to the generated mask. Discussing the performance, the no-fusion baseline achieves 27.11 dB PSNR and 0.818 SSIM. While MinMask fusion yields the lowest LPIPS (0.230), it severely degrades PSNR (26.90 dB) and SSIM (0.809). Median fusion improves over the baseline at 27.57 dB PSNR. Ultimately, our Softblend approach attains the best results at 27.70 dB PSNR and 0.829 SSIM. Consequently, we employ Softblend to combine multi-variant scene predictions in our final pipeline.